A kite is flying at a height of 23m and the string makes an angle of 53 degrees with the ground.


A) How long is the string (correct to 2 d.p)
B) If the kite is flying directly above a point X on the ground, how far is X from the person flying the kite (correct to the nearest m)?

Answer by TimothyLamb(4379)

---
sinA = opp/hyp
hyp = opp/sinA
hyp = length of string = 23/sin(53)
hyp = length of string = 28.8 m
---
tanA = opp/adj
adj = opp/tanA
adj = distance X from kite flyer = 23/tan(53)
adj = distance X from kite flyer = 17.3 m
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php